OpenStudy (firejay5):

Be Specific: How are circles and ellipses related? (Like at least 5 things)

OpenStudy (firejay5):

@campbell_st

OpenStudy (ikram002p):

#1 .the equation of ellips is \(x^2/a^2+ y^2/b^2 =r^2\) when a=b=c then \(x^2/c^2+ y^2/c^2 =r^2\) \(x^2 + y^2 =c^2r^2=s^2\) which is a circle . #2 . both are closed curve that also polar curves #3 you can say ellips about the circle , but you cant say circle bout the ellips #4 any point on ( circle or ellips ) have fixed sum distance with vocci  #5 both are semetric around x or y axis